OpenStudy (anonymous):

Find the inverse of g(x)=4x^3−15 I keep getting x^(1/3)((x+15)/4), what am I doing wrong?

OpenStudy (amistre64):

im going to assume everything ....

OpenStudy (amistre64):

y = 4x^3 -15 +15 +15 ---- ---- y+15 = 4x^3 /4 /4 y+15 ---- = x^3 ; cbrt the sides 4 y+15 cbrt (-----) = x 4

OpenStudy (amistre64):

then swap y for x since an inverse does that

OpenStudy (amistre64):

you actually got most of it good :)